# DreamAudio: Customized Text-to-Audio Generation with Diffusion Models

### v1

@@ IV-B3 Customized-Fantasy
We manually create a small-scale dataset as a benchmark for this new task, named as Customized-Fantasy. All the data clips are collected from online sources, including many unique and special events extracted from games or movies to simulate the user cases for real-world audio-production scenarios. For example, customization in laser guns, monster roars and Minion speaks. In addition, human interactions are involved to ensure that the captions are meaningful and present semantic coherence. In total, we collected 60 different events from Pixabay  [^1] https://pixabay.com/, a website which provides royalty-free audio resources under the Pixabay License suitable for research purposes, and compose $25$ distinct scenarios for zero-shot CTTA tasks.
@@ IV-B4 AudioCaps-General
To maintain the capability of the system in more robust situations, e.g., prompting on audio events without the reference concepts. We apply the general AudioCaps dataset with empty reference concepts. For instance, the model is given with the target audio clip and the target caption, along with three empty audio clips and empty text as the reference concept. The same training-testing split is applied for training and evaluation of the model.
@@ IV-C Evaluation Metrics
We use three different metrics for performance evaluation, including the Fréchet Audio Distance (FAD), Kullback-Leibler (KL) divergence and Contrastive Language-Pretraining (CLAP) score, by following the evaluation protocol of baseline audio generation models. In addition, we follow the evaluation pipelines from the customized image generations [22, 69] by incorporating subjective assessments that measure the consistency and the fidelity between generated and target audio, including CLAP audio score (CLAPA), AudioBox Aesthetics and subjective evaluations. Specifically, generative-based metrics including FAD, KL and CLAP-score are applied for evaluating the performance on Customized-Concatenation, Customized-Overlay and AudioCaps, subjective evaluations and AudioBox Aesthetics are used for the evaluations on Customized-Fantasy.
@@ IV-C2 KL Divergence
KL measures the logarithmic difference between the probabilities assigned by the distributions of two audio clips across all possible events. We apply the audio classification model PANNs [80] for the feature extraction and calculate the KL score between each target audio sample and their corresponding outputs. A lower KL presents a closer distance on distributions, suggesting the generated outputs are more similar to the target clips.
TABLE II: The setup of the primary experiments we performed, with DDPM for using DDPM-based training strategies and UNet for basic diffusion-based structure. DreamAudio with dual-encoder and single-encoder indicates whether the two groups of encoder blocks share the same weights.
@@
DreamAudio 279279M RFM Dual-Encoder
DreamAudio-L 891891M RFM Dual-Encoder
@@ IV-C3 CLAP
The CLAP score calculates the cosine similarity between the text embedding and the audio embedding provided by the CLAP model [51]. The CLAP model learns projectors to align the audio and text embeddings into a joint space and presents paired audio and language data with similar features within the latent space. Given both audio embedding ${e}_{a}$ and text embedding ${e}_{t}$, the score is calculated as:
$\text{CLAPScore}(x,r)=\frac{{e}_{a}\cdot{e}_{t}}{\text{max}(\|{e}_{a}\|\|{e}_{t}\|,\epsilon)},$} (7)
@@ IV-C4 CLAPA
Instead of calculating the similarity between audio feature and text feature, the CLAPA quantifies the relevance between the target audio feature embedding ${e}_{a}$ and the generated audio feature embedding $^{e}_{\hat{a}}$, by calculating their similarity as follows:
$\text{CLAPScore}_{A}(x,r)=\frac{{e}_{a}\cdot{e}_{\hat{a}}}{\text{max}(\|{e}_{a}\|\|{e}_{\hat{a}}\|,\epsilon)}.$} (8)
A higher score means that the generated output is more similar to the target audio within the feature space.
IV-C5 Subjective Evaluation
We adopt the evaluation metrics used in the baseline models [21], incorporating both Overall Impression (OVL) and Audio-Text Relation (REL) for subjective assessments. Detailed scoring instructions, including illustrative examples, are provided to the raters to ensure clarity. For OVL, we follow the approach outlined in the baseline models [21] by asking raters: How would you rate the overall quality of this audio sample?. Responses are scored on a Likert scale, ranging from 5 for “excellent” to 1 for “bad”. For REL, we adapt the metric to better suit the customized generation tasks by asking: Does the generated audio successfully present the target audio events customized by the reference audio samples? The subjective evaluations were performed on ten different listeners, with six researchers in audio and speech and four from other fields.
IV-C6 AudioBox Aesthetics
AudioBox Aesthetics is a recently proposed audio quality assessment tool that has demonstrated a strong correlation with human-based subjective evaluations [81]. In detail, this metric applies a transformer-based model for extracting and assessing the quality of audio clips without reference clips. In this paper, we choose the production quality (PQ) and content usefulness (CU) for the task. In detail, PQ is a technical objective aspect of the overall quality and CU is a subjective axis to evaluate the likelihood of audio sample for content creation.
@@ IV-C6 AudioBox Aesthetics
[…]
@@
DreamAudio AC+CM+CE $0.50$ $0.90$ $52.0$ $86.3$ $0.78$ $0.82$ $42.0$ $81.7$ $6.31$ $5.56$ $3.74$ $3.91$
DreamAudio-L AC+CM+CE $\mathbf{0.46}$ $\mathbf{0.92}$ $\mathbf{52.5}$ $\mathbf{87.7}$ $\mathbf{0.73}$ $\mathbf{0.67}$ $42.4$ $\mathbf{83.9}$ $\mathbf{6.37}$ $\mathbf{5.65}$ $\mathbf{3.89}$ $\mathbf{4.17}$
V Experiments Details
@@ V-A Model Architecture Details
For the text encoder, we use a pre-trained Flan-T5 model [53], resulting in a $1024$ dimension feature sequence with a fixed length of $50$ for every caption. We use the pre-trained VAE model from AudioLDM which provides a compression ratio of $4$ and results in a latent vector of $256$ dimension in temporal and $16$ dimension in frequency for a ten-second mel spectrogram. For the vocoder, we apply BigVGAN [50] and perform the self-supervised pre-training on $16$ kHz audio clips. Table II summarizes our experiments. We perform the experiments with two sizes of the latent diffusion model, DreamAudio and DreamAudio-L, with the hidden dimension size of $n_{\text{hidden}}=96$ and $n_{\text{hidden}}=128$, respectively. In addition, we performed evaluations of several variants of the proposed system. The DreamAudio system trained with DDPM strategies is denoted as DreamAudio-DDPM. The DreamAudio system with general UNet structure is named as DreamAudio-UNet. The system with shared-weight MRC structure is denoted as DreamAudio-SEncoder, where the encoder blocks for referenced feature extraction and feature encoding share the same weights.
@@ V-B Training and Inference Setup
Despite the pre-trained text encoder and VAE, we train the model and vocoder separately. We first train the vocoder on a large-scale audio dataset for 11M steps and then freeze the module into the proposed DreamAudio to train the model. Combining the whole system, we train the model with two customized datasets mentioned in Section IV-B. The system is trained on a single NVIDIA A100 80GB GPU for $2$ millions steps. Following similar strategies from generative networks, we apply the Classifier Free Guidance (CFG) for RFM with the value of $2.0$ for both DreamAudio and DreamAudio-L. We utilize the AdamW [82] optimizer with a learning rate of $5\times 10^{-5}$ for training and apply a linear warming up for $10,000$ steps.
@@ VI Results and Analysis
We evaluated our proposed system on both customized and general text-to-audio generation. The following section first discusses the performance of our system on the CTTA tasks and then explores the capability of general TTA tasks.
@@ VI-A Customized Text-to-Audio Generation
All the models in Table III are evaluated on both Customized-Concatenation and Customized-Overlay datasets. We compare our result with Re-AudioLDM [20] as this is the only work related to customized and zero-shot tasks. The result for AudioBox-TTA-RAG [29] is not included because the authors did not provide any model or checkpoint. For Re-AudioLDM, we apply the system with the retrieval number of $3$ as the retrieval-based information. As shown in Table III, the proposed DreamAudio significantly outperforms the previous systems across most of the metrics. The baseline system, Re-AudioLDM, achieves a FAD score of $3.05$ and $2.96$ for Customized-Concatenation and Customized-Overlay, respectively. Our systems, on the other hand, illustrate a substantially enhanced score of $0.46$ for the concatenated testing set and $0.73$ for the overlapped testing set. DreamAudio-L also attains the best KL divergence score of $0.92$ and $0.67$, considerably improved the performance upon any previous methods.
For the CLAP score, our system demonstrates SOTA performance on the Customized-Concatenation dataset and achieves comparable results on the Customized-Overlay dataset. Unlike other objective metrics that directly compare target and generated audio samples, the CLAP score measures semantic alignment rather than precisely evaluating the presence of customized audio content. In this case, it is less effective in evaluating the task that focuses on specific customization details. In contrast, the CLAPA score evaluates the similarity between the features from each generated audio clip and its corresponding target audio clip, making it a more accurate metric for reflecting the correctness of customized content. DreamAudio-L achieves a CLAPA score of $87.7$ on the Customized-Concatenation dataset and $83.9$ on the Customized-Overlay dataset, significantly surpassing all previous models. For the Customized-Fantasy testing set, DreamAudio achieves the best performance among production quality, content usefulness and human evaluation, where DreamAudio-L achieves SOTA results on OVL of $3.89$ and SOTA performance on REL of $4.17$. The significant enhancement offered by DreamAudio in both OVL and REL illustrates that the proposed system can generate specific audio features related to the reference concepts, as well as being faithful to text prompts.
TABLE IV: Comparison of model performances on the AudioCaps evaluation set. AC is short for AudioCaps dataset, CM is for Customized-Concatenation and CE is for Customized-Overlay. The dataset marked with $\ast$ indicates fine-tuning on AudioCaps.
@@ VI-B General Text-to-Audio Generation
In this section, we evaluate the capability of DreamAudio on general text-to-audio generation tasks. We compare the performance on audio generation with several SOTA systems, including AudioLDM [2], AudioGen [16], Make-an-Audio [19], Tango2 [52], AudioLDM2 [21], and Re-AudioLDM [20]. All the checkpoints are downloaded from open-sourced GitHub [^2] https://github.com/ or HuggingFace [^3] https://huggingface.co/ to reproduce the results. As shown in Table IV, these results demonstrate that our DreamAudio can maintain good performance in general text-to-audio tasks even without access to the corresponding customized content. For instance, DreamAudio fine-tuned on the AudioCaps dataset (providing empty reference contents) achieves a FAD score of $1.92$ and a KL score of $1.51$, which outperforms all non-retrieval methods. Compared to current SOTA models, DreamAudio achieves competitive CLAP scores of $47.5$ and $58.8$ in terms of semantic and audio feature similarity, respectively. DreamAudio did not outperform the baseline models probably because the MRC module did not provide useful audio information for general tasks. As a result, the model loses the advantage of incorporating external features and instead relies primarily on its own capability for independent audio feature generation.
Figure 5: The details of the MRC UNet network for reference length fine-tuning. All the previous blocks are frozen and only the new introduced CNN layer is trained during this stage.
TABLE V: The experiment on customized generation with four reference concepts, models are fine-tuned based the model with reference number of $3$, step number with indicates the model is directly applied without any further fine-tuning .
[…]
3 $0^{\ast}$ $0.50$ $92$ $52.5$ $87.7$
@@
VI-C Customized Generation on Various Reference Number
In this section, experiments are carried out to verify the capability of DreamAudio on various referenced concept pairs. To fit for different lengths of the referenced VAE latent, an extra CNN alignment layer is applied to squeeze the referenced feature so the concatenated feature being forwarded to the decoder blocks is designed as a fixed size, e.g. $96$ in the temporal dimension and $6$ in the frequency dimension. As depicted in Figure 5, we freeze all the blocks except the alignment blocks, which are trained when DreamAudio is fine-tuned for different lengths of reference concepts. The results of the proposed model for four reference pairs are presented in Table V. The experiment illustrates that DreamAudio can adapt the extended reference information within only $10,000$ steps and achieve a competitive FAD score of $0.79$ in only $50,000$ steps, which is comparable to the $0.50$ FAD score of DreamAudio on three pairs of reference concepts.
VI-D Ablation Studies
VI-D1 Effectiveness of RFM
As shown in Table III, we compare the outputs of DreamAudio with different training approaches on various datasets. Experiments on RFM-based systems present SOTA performance on customized tasks, i.e., Customized-Concatenation and Customized-Overlay evaluations. On the other hand, we observe that DreamAudio-DDPM trained with diffusion-based techniques offers better results in the semantic feature than DreamAudio, giving better CLAP scores. These experimental results demonstrate two key findings: a) RFM-based models show promising performance in generating customized and personalized content; b) DDPM-based approaches guide the models toward more robust performance by preserving the capability of generating features based on semantic information.
VI-D2 Effectiveness of MRC
We conduct various experiments with and without the UNet blocks with MRC in Table III For DreamAudio-UNet, the input dimension is extended four times to undertake both denoised latent and reference latent to align the process of MRC. The results illustrate that the proposed system with basic UNet structures exhibits significant performance degradation among all the metrics. In addition, the comparison between DreamAudio-SEncoder and DreamAudio indicates that the use of two groups of encoder blocks for encoding and feature extraction helps improve performance.
TABLE VI: Ablation Studies on the data masking, data dropping and the effectiveness of different proportions of training data. The general task is evaluated on the AudioCaps testing set and CTTA is evaluated on the Customized-Concatenation and Customized-Overlay testing sets. For data processing, AC is short for AudioCaps, and $C$ and CO are short for Customized-Concatenation and Customized-Overlay. The number indicates the quantity of each dataset for training.

### v2

@@ IV-B3 Customized-Fantasy
We manually created a small-scale dataset as a benchmark for this new task, named as Customized-Fantasy. All the data clips are collected from online sources, including many unique and special events extracted from games or movies to simulate the user cases for real-world audio-production scenarios. For example, customization in laser guns, monster roars and Minion speaks. In addition, human interactions are involved to ensure that the captions are meaningful and present semantic coherence. In total, we collected 60 different events from Pixabay [^1] https://pixabay.com/, a website which provides royalty-free audio resources under the Pixabay License suitable for research purposes. Together, we manually developed $25$ different scenarios to simulate real-world customized generation.
@@ IV-B4 AudioCaps-General
To maintain the capability of the system in more robust situations, e.g., prompting on audio events without the reference concepts. We apply the general AudioCaps dataset with empty reference concepts. For instance, the model is given with the target audio clip and the target caption, along with three empty waveform and empty text as the reference concept. The same training-testing split is applied for training and evaluation of the model.
@@ IV-C Evaluation Metrics
Following the evaluation protocol of baseline audio generation models [44], we use three different metrics for performance evaluation, including the Fréchet Audio Distance (FAD), Kullback-Leibler (KL) divergence and Contrastive Language-Pretraining (CLAP) related score. In addition, we follow the evaluation pipelines from the customized image generations [64, 13] and incorporate subjective assessments that measure the consistency and the fidelity between generated and target audio, including CLAP audio score (CLAPA), AudioBox Aesthetics, and subjective evaluations. Specifically, audio-text latent similarity based metrics including FAD, KL and CLAP-score are applied for evaluating the performance on Customized-Concatenation, Customized-Overlay and general AudioCaps testing sets, subjective evaluations and AudioBox Aesthetics are used for the evaluations on Customized-Fantasy.
@@ IV-C2 KL Divergence
KL measures the logarithmic difference between the probabilities assigned by the distributions of two audio clips across all possible events. We apply the audio classification model PANNs [34] for feature extraction and calculate the KL score between each target audio sample and their corresponding outputs. A lower KL represents a smaller distance in distributions, suggesting that the generated outputs are more similar to the target clips.
TABLE II: The setup of the primary experiments we performed, with DDPM for using DDPM-based training strategies, UNet for basic diffusion-based structure, and DiT for diffusion-transformer based backbone. DreamAudio with dual-encoder and single-encoder indicates whether the two groups of encoder blocks share the same weights.
@@
DreamAudio-DiT-B 920920M RFM Dual-Encoder
DreamAudio-DiT-L 1.31.3B RFM Dual-Encoder
@@
DreamAudio 891891M RFM Dual-Encoder
DreamAudio-L 1.11.1B RFM Dual-Encoder
@@ IV-C3 CLAP
The CLAP score calculates the cosine similarity between the text embedding and the audio embedding provided by the CLAP model [84]. The CLAP model learns projectors to align the audio and text embeddings into a joint space and presents paired audio and language data with similar features within the latent space. Given both generated audio embedding $\mathbf{e}_{a}$ and text embedding $\mathbf{e}_{t}$, the score is calculated as:
$\text{CLAP}(\mathbf{e}_{a},\mathbf{e}_{t})=\frac{\mathbf{e}_{a}\cdot\mathbf{e}_{t}}{\text{max}(\|\mathbf{e}_{a}\|\|\mathbf{e}_{t}\|,\epsilon)},$} (9)
@@ IV-C4 CLAPA
Instead of calculating the similarity between audio and text embeddings, the CLAPA score is specially designed for the CTTA task to compare the customized concepts between the targets and generated audio by calculating the cosine similarity score between the target audio embedding $\mathbf{e}_{a}$ and the generated audio embedding $\mathbf{e}_{\hat{a}}$,
$\text{CLAP}_{A}(\mathbf{e}_{a},\mathbf{e}_{\hat{a}})=\frac{\mathbf{e}_{a}\cdot\mathbf{e}_{\hat{a}}}{\text{max}(\|\mathbf{e}_{a}\|\|\mathbf{e}_{\hat{a}}\|,\epsilon)}.$} (10)
A higher score means that the generated output is more similar to the target audio within the embedding space.
IV-C5 AudioBox Aesthetics
AudioBox Aesthetics is a recently proposed audio quality assessment tool that has demonstrated a strong correlation with subjective evaluations by human listeners [79]. In detail, this metric applies a transformer-based model for extracting and assessing the quality of audio clips without reference clips. In this paper, we choose the production quality (PQ) and content usefulness (CU) for the task. In detail, PQ is an objective aspect of the overall quality and CU is a subjective axis to evaluate the likelihood of audio sample for content creation.
IV-C6 Subjective Evaluation
We adopt the evaluation metrics used in the baseline models [44], incorporating both Overall Impression (OVL) and Audio-Text Relation (REL) for subjective assessments. Detailed scoring instructions, including illustrative examples, are provided to the raters to ensure clarity. For OVL, we follow the approach outlined in the baseline models [44] by asking raters: How would you rate the overall quality of this audio sample?. Responses are scored on a Likert scale, ranging from 5 for “excellent” to 1 for “bad”. For REL, we adapt the metric to better suit the customized generation tasks by asking: Does the generated audio successfully present the target audio events customized by the reference audio samples? The subjective evaluation was conducted with ten human raters recruited through an open advertisement within the Department of Electrical and Electronic Engineering at the University of Surrey. All participants were PhD students and were not involved in the development of the proposed model or the writing of this paper, and none of the authors participated in the evaluation. To encourage diverse perspectives and reduce potential bias, the participant pool included six raters with backgrounds in audio related research and four raters from unrelated fields such as computer vision and robotics, providing complementary non-expert assessments.
@@ IV-C6 Subjective Evaluation
FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ CLAPA $\uparrow$ FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ CLAPA $\uparrow$ PQ $\uparrow$ CU $\uparrow$ OVL $\uparrow$ REL $\uparrow$
@@
DreamAudio AC+CM+CE $0.50$ $\mathbf{0.90}$ $52.0$ $86.3$ $0.78$ $0.82$ $42.0$ $81.7$ $6.31$ $5.56$ $3.74$ $3.91$
DreamAudio-L AC+CM+CE $\mathbf{0.46}$ $0.92$ $\mathbf{52.5}$ $\mathbf{87.7}$ $\mathbf{0.73}$ $\mathbf{0.67}$ $42.4$ $\mathbf{83.9}$ $6.37$ $5.65$ $\mathbf{3.89}$ $\mathbf{4.17}$
V Experimental Setting
@@ V-A Model Architecture Details
For the text encoder, we use a pre-trained Flan-T5 model [60], resulting in a $1024$ dimension feature sequence with a fixed length of $50$ for every caption. We use the pre-trained VAE model from AudioLDM [43] which provides a compression ratio of $4$ and results in a latent vector of $256$ dimension in temporal and $16$ dimension in frequency for a ten-second mel spectrogram. For the vocoder, we apply BigVGAN [42] and perform the self-supervised pre-training on audio clips sampled at $16$ kHz. Table II summarizes our experiments. We perform the experiments with two sizes of the latent diffusion model, DreamAudio and DreamAudio-L, with the hidden dimension size of $n_{\text{hidden}}=96$ and $n_{\text{hidden}}=128$, respectively. In addition, we evaluated several variants of the proposed system. The DreamAudio trained under a DDPM framework is denoted as DreamAudio-DDPM. The version using a standard UNet backbone but without the MRC module is named DreamAudio-UNet. In this system, rather than using the dual-path MRC structure, all three reference audio latent features are concatenated together with the latent of the noisy input at the input layer, extending the input dimension to align with the MRC process. The system with shared-weight MRC encoders is denoted as DreamAudio-SEncoder, where the encoder blocks used for referenced feature extraction and target feature encoding share the same encoder. This encoder is simply run multiple times for the target latent and each referenced latent during training and inference. Beyond the UNet-based architectures, we also experimented with two models that adopt a diffusion-transformer (DiT) backbone [55]. In DreamAudio-DiT-B, we employ 12 transformer layers, while in DreamAudio-DiT-L, we use 24 transformer layers. Both models incorporate skip connections, and first half of the layers are used as the encoders while the second half is applied as the decoders.
@@ V-B Training and Inference Setup
Despite the pre-trained text encoder and VAE, we train the model and vocoder separately. We first train the vocoder on a large-scale audio dataset for 11M steps and then freeze the module while we train the proposed DreamAudio. We collected $288,283$ clips for model training and $1328$ clips for model evaluation by combining Customized-Concatenation, Customized-Overlay, and general AudioCaps datasets mentioned in Section IV-B. For each audio clip, we provide three-pairs of reference concepts as the external data. The system is trained on a single NVIDIA A100 80GB GPU for $2$ millions steps. Following similar strategies from generative networks, we apply the Classifier Free Guidance (CFG) for RFM with a value of $2.0$ for both DreamAudio and DreamAudio-L. We utilize the AdamW [52] optimizer with a learning rate of $5\times 10^{-5}$ for training and apply a linear warming up for $10,000$ steps.
@@ VI Results and Analysis
We evaluated DreamAudio on both customized and general text-to-audio generation. The following section first discusses the performance of our system on the CTTA tasks and then explores the capability of general TTA tasks.
@@ VI-A Customized Text-to-Audio Generation
All the models in Table III are evaluated on both Customized-Concatenation and Customized-Overlay datasets. We compare our method with Re-AudioLDM [91] as this is the only work related to customized and zero-shot tasks. The result for AudioBox-TTA-RAG [86] is not included because the authors did not provide any model or checkpoint. For Re-AudioLDM, we apply the system with the retrieval number of $3$ as the retrieval-based information. As shown in Table III, the proposed DreamAudio significantly outperforms the previous systems across most of the metrics. The baseline system, Re-AudioLDM, achieves a FAD score of $3.05$ and $2.96$ for Customized-Concatenation and Customized-Overlay, respectively. Our systems, on the other hand, illustrate a substantially enhanced score of $0.46$ for the concatenated testing set and $0.73$ for the overlapped testing set. DreamAudio-L also achieves the best KL divergence score of $0.92$ and $0.67$ on the two datasets, respectively.
For the CLAP score, our system demonstrates the best performance on the Customized-Concatenation dataset and achieves comparable results on the Customized-Overlay dataset. Unlike other objective metrics that directly compare target and generated audio samples, the CLAP score measures semantic alignment rather than precisely evaluating the presence of customized audio content. In this case, it is less effective in evaluating the CTTA task. In contrast, the CLAPA score evaluates the similarity between the features from each generated audio clip and its corresponding target audio clip, making it a more accurate metric for reflecting the correctness of customized content. DreamAudio-L achieves a CLAPA score of $87.7$ on the Customized-Concatenation dataset and $83.9$ on the Customized-Overlay dataset, significantly surpassing all previous models. For the Customized-Fantasy testing set, DreamAudio achieves the best performance among production quality, content usefulness and human evaluation, where DreamAudio-L achieves better results on OVL of $3.89$ and REL of $4.17$. The significant enhancement of these metrics illustrates that DreamAudio can generate specific audio features related to the reference concepts and be faithful to text prompts.
TABLE IV: Comparison of model performances on the AudioCaps evaluation set. AC is short for AudioCaps dataset, $C$ is for Customized-Concatenation and CO is for Customized-Overlay. The dataset marked with $\ast$ indicates fine-tuning and means training from scratch on AudioCaps.
@@ VI-B General Text-to-Audio Generation
In this section, we evaluate the capability of DreamAudio on general text-to-audio generation tasks. We compare the performance on audio generation with several SOTA systems, including AudioLDM [43], AudioGen [38], Make-an-Audio [27], Stable Audio Open [16], Tango2 [22], AudioLDM2 [44], TangeFlux [28], and Re-AudioLDM [91]. All the checkpoints are downloaded from open-sourced GitHub or HuggingFace to reproduce the results. For plain text-to-audio tasks without any reference inputs, DreamAudio is provided with empty reference audio clips and empty reference text token. As shown in Table IV, DreamAudio trained on reference-based data does not achieve SOTA performance when directly applied to general text-to-audio generation tasks. We believe the main reason for this is that DreamAudio is architecturally designed for customized generation with reference-based conditions. The MRC module provides an additional pathway for extracting and leveraging external audio/text reference signals. When reference data are available during training stages, the model learns to rely on these external concepts for capturing acoustic feature and fine-grained attributes. This makes the model less dependent on the text conditions while generating audio. Consequently, when operating on general tasks without reference concepts, the external pathway of MRC is un-used. Hence, the performance is limited since the model has been guided by the conditions from both reference concepts and text prompts, rather than purely text prompts.
After fine-tuning DreamAudio on AudioCaps, the model recovers strong performance and achieves comparable results to most non-retrieval based systems. Furthermore, when DreamAudio is trained from scratch with only AudioCaps and no reference-based data, the model exhibits even better performance, demonstrating that the architecture itself remains fully capable of high-quality general text-to-audio generation. These findings indicate that the system can be applied to general text-to-audio generation tasks by reducing the conditioning of the reference concepts.
Figure 5: The details of the MRC UNet network for reference length fine-tuning. All the existing blocks are frozen and only the introduced CNN alignment layer is trained during this stage.
TABLE V: Experimental results for customized generation with four reference concepts, where the models are obtained by fine-tuning the model trained with 33 references. The step number with indicates that the model is fully-trained with 33 references and without any further fine-tuning.
Reference Num Steps FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ CLAPA $\uparrow$
3 $0^{\ast}$ $0.50$ $0.92$ $52.5$ $87.7$
@@
VI-C Customized Generation on Various Number of References
Our DreamAudio is primarily trained and evaluated with three-pairs of reference concepts. In this section, we investigate the capability of DreamAudio to perform customized generation with four pairs of reference concepts. The current backbone adopts a CNN-based U-Net architecture, where all intermediate latent feature maps within the decoder must maintain fixed spatial dimensions to allow their concatenation through skip connections. In this case, the referenced VAE latent features need to be projected into a fixed-size representation, e.g. $96$ in the temporal dimension and $6$ in the frequency dimension for three reference pairs. When the system is given more than three-pairs of reference concepts, the dimension of reference latent feature provided by the feature extraction path does not match with the designed input. To address this issue, we introduce a lightweight CNN alignment layer that maps the reference VAE inputs with variable temporal lengths into this fixed-size latent space, as shown in Figure 5.
During fine-tuning, all backbone encoder–decoder blocks are frozen and only the CNN alignment layer is trained to adapt to the scenarios with different numbers of reference concepts. The performance results of DreamAudio trained or fine-tuned with four-pairs referenced concepts are shown in Table V. The first row shows the performance of a fully-trained DreamAudio model using 3 reference concepts. In contrast, a model using 4 references without any fine-tuning (0 steps) fails, as its untrained CNN alignment layer cannot properly reshape or compress the four latent inputs into a compatible, fixed-size representation. This results in meaningless outputs and a sharp drop in performance. However, after just 10,000 fine-tuning steps, the alignment layer learns to effectively project the four references into a unified latent space, adapting quickly to the additional input. Performance continues to improve with extended training, achieving a competitive FAD score of 0.79 after 50,000 steps. These results demonstrate that DreamAudio can readily generalize to different numbers of reference concepts by integrating a simple, trainable CNN alignment layer.
TABLE VI: Ablation studies on data augmentation (masking, dropping) and the effectiveness of different proportions of training data. The general task is evaluated on the AudioCaps testing set and CTTA is evaluated on the Customized-Concatenation and Customized-Overlay testing sets. For data processing, AC is short for AudioCaps, and $C$ and CO are short for Customized-Concatenation and Customized-Overlay. The number indicates the quantity of each dataset for training. The result with indicates the experiments without reference text concepts.